a room is 15 feet by 30 feet and the tile 1 foot and cost 92 cent how much will it cost to cover the entire room
Answer by mananth(16949)

area of room /area of tile = number of tiles
number of tiles * cost per tile = Total cost of tiling.
15*30 = 450 sq.ft
tile area = 1*1=1 sq.ft
450/1 = 450 tiles
450 *0.92 = $414